Groups of campers where going to an island on the first day 10 went over and to come back on the second day 12 went over and third three came back is this pattern continuous how many would be on this island at the end of the week? And how many would have left ?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Groups of campers were going to an island, On the first day 10 went over and to came back. On the second day 12 went over and three came back. Is this pattern continuous? How many would be on this island at the end of the week? And how many would have left ?
---------
1st day:: 10-2 = 8 on island
2nd day:: 12-3 = 9 + 8 on = 17 on island
----
Making some assumptions:
3rd day:: 14-4 = 10 + 17 = 27 on island
4th day:: 16-5 = 11 + 27 = 38 on island
etc
------
Your post is not clear enough to determine the pattern for
days 3, 4, 5 etc.
--------------
Cheers,
Stan H.
-----------

Answer by ikleyn(52781)

.
WHICH PATTERN ?

There is no any pattern in this post.